Brad is eight more than two times as old as Sarah. In 10 years the sum of their ages withh be 76. How old is Sarah now?
Answer by stanbon(75887)

Brad is eight more than two times as old as Sarah. In 10 years the sum of their ages withh be 76. How old is Sarah now?
---------
Equations:
b = 2s + 8
b+10 + s + 10 = 76
-----
Substitute for "b" and solve for "s":
2s+8 + 10 + s+10 = 76
3s + 28 = 76
3s = 48
s = 16 yrs (Sarah's age now)
----
Solve for "b":
b = 2s + 8
Brad = 40 yrs (Brad's age now)
----------
Cheers,
Stan H.
=============